11) (2 pt) pH relies upon the dissociation of H2O to H+ and OH- . In a bacterium with a volume of 10-15 L, how many hydronium ions are present? If the inner mitochondrial matrix is at a volume of 0.9 femtoliter and pH 8.05, how many hydronium ions are present? Simplistically, if 3 H+ must be pumped out of the matrix for each ATP formed, how many ATP could be produced?
Answer #1

Solution :-

Using the pH we can find the concentration of the hydronium ion.

pH=-log[H^+]

[H^+] =10^(-pH)
           =10^(-8.05)
           =8.91*10^-9 M

Now using the molarity and volume we can find the moles of H^+

Moles= molarity x volume in L

Converting 0.9 femtoliter to liter

0.9 femtoliter *(1L/1*10^-15 femtoliter) =9*10^-16 L

Calculating moles of H^+

Moles of H^+ = 8.91*10^-9 mol per L *9*10^-16 L
                          =8.02*10^-24 mol H^+

Converting moles to number of H^+ ions

(8.02*10^-24 mol H^+)*(6.022*10^23 / 1 mol )=4.83 H^+

So number of hydronium ions present are 4.83 H^+

Now calculating the ATP

(4.83 H^+)*(1 ATP /3 H^+) =1.61 ATP

So it can produce 1.61 ATP